# Yong Liu    Guo Qin    Zhiyuan Shi    Zhi Chen    Caiyin Yang    Xiangdong Huang    Jianmin Wang    Mingsheng Long

### v1

@@ Abstract
We introduce Sundial, a family of native, flexible, and scalable time series foundation models. To predict the next-patch’s distribution, we propose a TimeFlow Loss based on flow-matching, which facilitates native pre-training of Transformers on time series without discrete tokenization. Conditioned on arbitrary-length time series, our model is pre-trained without specifying any prior distribution and can generate multiple probable predictions, achieving flexibility in representation learning beyond using parametric densities. Towards time series foundation models, we leverage minimal but crucial adaptations of Transformers and curate TimeBench with 1 trillion time points, comprising mostly real-world datasets and synthetic data. By mitigating mode collapse through TimeFlow Loss, we pre-train a family of Sundial models on TimeBench, which exhibit unprecedented model capacity and generalization performance on zero-shot forecasting. In addition to presenting good scaling behavior, Sundial achieves new state-of-the-art on both point forecasting and probabilistic forecasting benchmarks. We believe that Sundial’s pioneering generative paradigm will facilitate a wide variety of forecasting scenarios.
@@ 1 Introduction
Time series forecasting has fascinated people for thousands of years. Although people have been able to determine the time using instruments like sundials in 3000 BC, time series forecasting is intrinsically non-deterministic (Box et al., 2015). Therefore, generating the range of probable predictions is crucial for decision-making. The growing demand has facilitated numerous statistical approaches over the past decades (Hyndman, 2018; Box, 2013), which provide high-profile theories and probabilistic models to make reliable schedules. Recent advancements bring the boom of deftly designed models that automatically learn intricate dynamics and correlations from raw data (Oreshkin et al., 2019; Nie et al., 2022; Zhang & Yan, 2023; Liu et al., 2023). Despite the impressive performance, deep models necessitate task-specific training on sufficient in-distribution data. Motivated by advances in large models (Bommasani et al., 2021), pre-trained time series foundation models have shown promising capabilities in out-of-distribution tasks (Das et al., 2023b; Liu et al., 2024b; Woo et al., 2024; Ansari et al., 2024).
@@
Figure 1: A native time series model learns the representation on the continuous token. A flexible foundation model is pre-trained without specifying the prior distribution. Sundial is presented as the first family of native and flexible time series foundation models.
@@
Current research of time series foundation models has converged on building unified, scalable, and out-of-the-box forecasters, exhibiting zero-shot performance close to or sometimes surpassing supervised methods (Aksu et al., 2024). Notably, Transformers (Radford et al., 2018) are currently the de facto architecture of these models. While pre-trained Transformers with inherent generative capability have facilitated great success in language, image, and video generation (Ramesh et al., 2021; OpenAI, 2023; Liu et al., 2024c), most time series foundation models are not “generative” or, more specifically, probabilistic forecasters, thereby limiting reliability in decision-making. Although parametric densities specified with prior distributions (Wen et al., 2017; Woo et al., 2024) can be incorporated to cope with the uncertainty in predictions, they can constrain the capacity of distributions represented by a foundation model, especially on time series corpora characterized by high heterogeneity. To learn arbitrarily-intricate distributions without mode collapse, language modeling (Bengio et al., 2000) that learns the categorical distribution via cross-entropy loss inspires subsequent works (Gruver et al., 2023; Ansari et al., 2024), which treat time series as a foreign language using discrete tokenization. Still, apparent distinctions between continuous-valued time series and language tokens can lead to out-of-vocabulary issues and coarse-grained prediction intervals.
@@
As shown in Figure 1, we leverage the native patching for time series tokenization and generative modeling to learn flexible predictive distributions for the first time. As foundation models aim to learn complicated distributions from extensive datasets and facilitate transferability across agnostic downstream tasks, we do not adopt any specific probabilistic priors, such as unimodal Gaussian or multimodal mixtures. To tame Transformers as native and scalable time series foundation models, we adopt generative modeling, a well-established paradigm that offers as much flexibility as language modeling and is better suited for continuous-value time series. Instead of adopting prevailing denoising diffusion models (Ho et al., 2020), we adopt a simple yet effective flow-matching framework (Lipman et al., 2022), which provides better efficiency and quality for generation (Tong et al., 2023). We propose TimeFlow Loss, formulated as a parameterized training objective, for autoregressive models to learn each token’s predictive distribution. This optimization objective accurately operates on original values and facilitates patch-level generation for quick inference, which is highly compatible with continuous-valued modalities.
@@
Besides the TimeFlow Loss, we enhance Transformers with minimal but critical adaptations. We design patch embedding that is compatible with non-divisible context length. We adopt RoPE (Su et al., 2024) to enhance temporal causality. We leverage Pre-LN (Xiong et al., 2020), FlashAttention (Dao et al., 2022), and KV Cache (Pope et al., 2023), which are crucial but generally neglected in the development of time series foundation models. With the TimeFlow Loss facilitating the right paradigm for training scalable foundation models, these adaptations further optimize deployment.
@@
To explore the scaling law of time series foundation models, we collect and curate TimeBench with an unprecedented volume of 1 trillion time points. We present Sundial as a family of highly capable foundation models, which achieve state-of-the-art on three large-scale and best-recognized benchmarks, including Time-Series-Library (TSLib) (Wu et al., 2022), GIFT-Eval (Aksu et al., 2024) and FEV (Ansari et al., 2024). Our contributions lie in these aspects:
@@
  • We present Sundial, a family of scalable and efficient time series foundation models pre-trained on 1 trillion time points, utilizing our enhanced Transformer.
  • Experimentally, Sundial achieves state-of-the-art zero-shot performance on point forecasting benchmarks and probabilistic forecasting leaderboards, including GIFT-Eval and FEV, positioning generative time series foundation models as a capable tool for decision-making.
@@ 2.1 Time Series Forecasting
Forecasting is essential for decision-making, which has facilitated the development of statistical and deep-learning models. Advancements in deep models for time series include theory-inspired components (Wu et al., 2021; Zeng et al., 2023; Wu et al., 2022), architecture-oriented adaptations (Bai et al., 2018; Salinas et al., 2020; Lim et al., 2021), and time series processing (Kim et al., 2021; Nie et al., 2022). While deep models learning the dataset-level distribution enjoy large model capacity, statistical methods fitting separate dynamics of each time series are still prevailing choices due to their flexibility and performance on small data (Ke et al., 2017; Hyndman, 2018).
One of the efforts towards more capable forecasters focuses on the foundation models (Bommasani et al., 2021), which address data-scarce scenarios by pre-training. Recent models aim to generalize on out-of-distribution data, achieving training-free overheads like statistical methods while enjoying the capacity as deep models. Another aspect is the shift from point to probabilistic forecasting (Woo et al., 2024; Ansari et al., 2024), which greatly improves the experience across many use cases by addressing forecasting uncertainty. While parametric densities can be easily incorporated into task-specific training, they can be overwhelmed by the heterogeneity of large-scale corpora. Applying this paradigm to pre-train foundation models is likely to result in mode collapse, manifested as over-smooth predictions given by the pre-trained model. In this work, we formally introduce generative time series foundation models, which naturally address the uncertainty in forecasting.
@@ 2.2 Time Series Foundation Models
Recent research has concentrated on building versatile large time series models (Liang et al., 2024). With the advance made in large language models, Transformer has become the dominant architecture. Several works adapt Transformers to address the unique 2D-dimensionality and heterogeneity of time series (Woo et al., 2024; Liu et al., 2024a). Specifically, our work delves into tokenization and optimization. Models such as TimesFM (Das et al., 2023b), Timer (Liu et al., 2024a, b), and Time-MoE (Shi et al., 2024b) embed continuous values and fit unimodal distributions via MSE or quantile loss (Wen et al., 2017). However, such prior loss may result in mode collapse in the of pre-training foundation models, and deterministic outcomes often fail to satisfy the requirement of decision-making. Based on continuous tokenization, Moirai (Woo et al., 2024) is a probabilistic model learning a mixture of distributions, but this prior can still struggle to fit complex distributions. Inspired by language modeling for scalable pre-training, Chronos (Ansari et al., 2024) discretizes series by scaling and quantization, learning more flexible categorical distribution by cross-entropy. Still, discrete tokenizer is limited to point level and sensitive to quantization interval, which is replaced by patch embedding and quantile head in subsequent work. Unlike before, we tame Transformer as native time series foundation model, learning flexible distributions without discrete tokenization.
@@ 2.3 Generative Modeling for Time Series
To address distributional heterogeneity during pre-training, generative modeling has become a focal point in the development of foundation models (Zhao et al., 2023; Liu et al., 2024c). While this direction for time series mostly focused on time series generation (Tashiro et al., 2021) and task-specific forecasters (Rasul et al., 2021; Shen & Kwok, 2023; Kollovieh et al., 2024), generative modeling for time series foundation models is hardly explored. With the comparable flexibility in distribution learning as language modeling, diffusion denoising (Sohl-Dickstein et al., 2015) and flow-matching (Lipman et al., 2022) have gained increasing prevalence in continuous-valued modalities (Lipman et al., 2024). Compared with diffusion denoising models, flow-matching provides a simple yet efficient framework. With fewer steps involved in the forward and reverse processes, large models based on flow-matching have shown superior performance in image generation (Esser et al., 2024).
Despite the connection in value continuity, generating images and future time series are fundamentally different tasks due to the autoregressive property of forecasting. Our proposed TimeFlow Loss is designed for autoregressive models to conduct conditional generation, which is a parameterized loss function (Zhang et al., 2018) for arbitrary distributions that enhances representation learning of foundation models.
@@ 3.2 Generative Models for Probabilistic Forecasting
Given a historical observation $x_{1:t}=\{x_{1},...,x_{t}\}$, the target of time series forecasting is to predict future time series $x_{t+1:t+f}=\{x_{t+1},\dots,x_{t+f}\}$. The task can be generally formulated as $p(x_{t+1:t+f}|\mathbf{h}_{t})$, where $\mathbf{h}_{t}=f_{\phi}(x_{1:t})$ is the learned representation from a deep model $f_{\phi}$. In probabilistic forecasting, explicit optimization objectives are utilized to predict the statistics of future series, e.g., MSE or quantiles, which have specified $p$ as a prior distribution. While using one parametric density generally fits well on a small amount of data, it can be the major bottleneck for scaling time series foundation models. Inspired by the success of large generative models (Rombach et al., 2022; OpenAI, 2023; Esser et al., 2024), we introduce generative modeling to realize probabilistic forecasting:
@@
$g_{\theta}$ is a small trainable generative network conditioned on the learned representations of $f_{\phi}$. Both of them are jointly optimized. While the generative model automatically learns the distribution, it supports probabilistic forecasting by generating raw predictions and calculating their statistics. The idea is conceptually related to conformal prediction (Angelopoulos & Bates, 2021) but enjoys flexibility in outputs.
@@ 4 Approach
In this work, we conduct a univariate pre-training paradigm, which adopts Channel-Independence (Nie et al., 2022) on multivariate data. To mitigate value range discrepancy in time series, we conduct normalization on time series individually per variable. Afterwards, we sample varying-length training samples with the maximum context length of $2880$. As a foundation model, Sundial is required to predict on out-of-distribution series with varied lengths during inference.
@@ 4.1 Sundial
As shown in Figure 2, our Sundial models consist of three parts: (1) time series tokenization, including a context-level re-normalization and a patch embedding that addresses any-length time series, (2) a Transformer backbone that learns the per-token representation of time series, and (3) TimeFlow Loss, a parameterized loss function to model the per-token distribution and generate raw series during inference. Sundial is designed to operate on continuous-valued time series and facilitates scalable pre-training on heterogeneous distribution. Additionally, Sundial presents a new approach to probabilistic forecasting via generative modeling.
@@ 4.1.1 Time Series Tokenization
We adopt re-normalization (Liu et al., 2022), a non-parametric two-stage instance normalization conducted within each sample. While it is initially proposed to mitigate non-stationarity of time series, it addresses the value discrepancies and temporal distribution shifts, improving generalizability for zero-shot forecasting.
@@
where $\mathbf{h}_{i}\in\mathbb{R}^{D}$ and $D$ is the dimension of token embedding. Unlike point-level quantization (Ansari et al., 2024), we reserve original values without discrete quantization while reducing the number of tokens feeding to the Transformer.
@@ 4.1.2 Transformer Backbone
Given $N$ token embeddings $\{\mathbf{h}_{i}\}$, we adopt three crucial adaptations on a decoder-only Transformer to obtain per-token representations aggregated from all previous tokens. First, we adapt Pre-LN (Xiong et al., 2020) to improve pre-training stability. Second, we leverage a causal self-attention mechanism with RoPE (Su et al., 2024) that introduces the position information of patch tokens. It can be formulated as follows (the layer index is omitted for simplicity):

### v2

@@ Abstract
We introduce Sundial, a family of native, flexible, and scalable time series foundation models. To predict the next-patch’s distribution, we propose a TimeFlow Loss based on flow-matching, which facilitates native pre-training of Transformers on continuous-valued time series without discrete tokenization. Conditioned on arbitrary-length time series, our models are pre-trained without specifying any prior distribution and can generate multiple probable predictions, achieving more flexibility in representation learning than using parametric densities. Towards time series foundation models, we leverage minimal but crucial adaptations of Transformers and curate TimeBench with one trillion time points, comprising mostly real-world datasets and synthetic data. By mitigating mode collapse via TimeFlow Loss, we pre-train a family of Sundial models on TimeBench, which achieve unprecedented model capacity and generalization performance. In addition to excellent scalability, Sundial achieves state-of-the-art results on both point and probabilistic forecasting benchmarks with a just-in-time inference speed, i.e., making zero-shot predictions within a few milliseconds. We believe that Sundial’s pioneering generative forecasting capability can improve model reliability in real-world decision-making. Code is available at: https://github.com/thuml/Sundial.
@@ 1 Introduction
Time series forecasting has fascinated people for thousands of years. Although people have been able to determine the time using instruments like sundials in 3000 BC, time series forecasting is intrinsically non-deterministic (Box et al., 2015). Therefore, generating a variety of probable predictions is crucial for decision-making. The growing demand has facilitated numerous statistical approaches over the past decades (Hyndman, 2018; Box, 2013), which provide high-profile theories and probabilistic tools for making reliable schedules. Recent advancements bring the boom of deftly designed models that automatically learn intricate dynamics and correlations from raw data (Oreshkin et al., 2019; Nie et al., 2022; Zhang & Yan, 2023; Liu et al., 2023a). Despite the impressive performance, deep models necessitate task-specific training on sufficient in-distribution data. Motivated by advances in large models (Bommasani et al., 2021), pre-trained time series foundation models have shown promising capabilities in out-of-distribution tasks (Das et al., 2023b; Liu et al., 2024b; Woo et al., 2024; Ansari et al., 2024).
@@
Figure 1: A native time series model operates on the original series of continuous values. A flexible foundation model is pre-trained without specifying prior distributions. Sundial is the first family of native and flexible time series foundation models.
@@
Current research on time series foundation models has converged on building unified, scalable, and out-of-the-box forecasters, exhibiting zero-shot performance close to or sometimes surpassing supervised methods (Aksu et al., 2024). Notably, Transformers (Radford et al., 2018) are currently the de facto architecture of these models. While pre-trained Transformers with an inherent generative ability have facilitated great success in language, image, and video generation (Ramesh et al., 2021; OpenAI, 2023; Liu et al., 2024c), most time series foundation models are not “generative” or, more specifically, probabilistic forecasters, thereby limiting reliability in decision-making. Although parametric densities specified with prior distributions (Wen et al., 2017; Woo et al., 2024) can be adopted to address uncertainty in time series forecasting, they can reduce the capacity of distributions learned by pre-trained models, especially on time series modality characterized by high heterogeneity. To learn arbitrarily intricate distributions without mode collapse, language modeling (Bengio et al., 2000) that learns the categorical distribution via cross-entropy loss inspires subsequent works (Gruver et al., 2023; Ansari et al., 2024), which treat time series as a foreign language using discrete tokenization. Still, discrepancies between continuous-valued time series and discrete language tokens can lead to out-of-vocabulary issues and coarse-grained prediction intervals.
@@
As shown in Figure 1, Sundial is presented as the first family of generative models among time series foundation models. As foundation models intend to learn complicated distributions from extensive datasets and facilitate transferability across agnostic downstream datasets, we do not specify any prior parametric densities, such as unimodal and multimodal Gaussian mixtures. Instead, we delve into generative modeling to tame Transformers as native, flexible, and scalable time series foundation models. By comparing to denoising diffusion models (Li et al., 2024), we opt for a simple yet effective flow-matching framework (Lipman et al., 2022), which provides notable efficiency and sample quality (Tong et al., 2023). We propose TimeFlow Loss, a parameterized training objective (Zhang et al., 2018) for autoregressive models to learn and sample from each token’s predictive distribution. Optimizing models in the original continuous-valued domain, TimeFlow Loss facilitates patch-level generation and enables fast inference, which is naturally compatible with the time series modality.
@@
In addition to TimeFlow, we enhance the Transformer with minimal but critical adaptations. We develop feasible patch tokenization for arbitrary-length input time series. We adopt RoPE (Su et al., 2024), Pre-LN (Xiong et al., 2020), FlashAttention (Dao et al., 2022), and KV Cache (Pope et al., 2023), which are crucial but generally neglected in the development of time series foundation models. Besides, we pre-train our models by multi-patch prediction to reduce autoregression steps. We realize a rapid generation of multiple samples by reusing a shared lookback representation. Beyond facilitating scalable pre-training, these adaptations help real-time long-context inference and long-term generation.
@@
To validate the scaling law of time series foundation models, we collect and curate TimeBench with an unprecedented volume of a trillion time points. We present Sundial as a family of highly capable foundation models, which achieve state-of-the-art on three large-scale and best-recognized benchmarks, including Time-Series-Library (TSLib) (Wu et al., 2022), GIFT-Eval (Aksu et al., 2024), and FEV (Ansari et al., 2024). Our contributions lie in these aspects:
@@
  • We present Sundial, a family of scalable and efficient time series foundation models built upon our enhanced Transformer and pre-trained on a trillion time points.
  • Experimentally, Sundial achieves state-of-the-art zero-shot performance on point and probabilistic forecasting benchmarks, including TSLib, GIFT-Eval, and FEV, indicating a promising generative approach for the future improvement of time series foundation models.
@@ 2.1 Time Series Forecasting
Forecasting is essential for decision-making. Advancements in deep learning for time series include theory-inspired deep modules (Wu et al., 2021; Liu et al., 2023b; Wu et al., 2022), architecture-oriented adaptations (Bai et al., 2018; Salinas et al., 2020; Lim et al., 2021), and time series preprocessing (Kim et al., 2021; Nie et al., 2022). Deep models learn the dataset-level distribution and benefit from strong generalization and model capacity. Statistical methods conduct case-by-case fitting on input series, achieving notable performance on small data (Ke et al., 2017; Hyndman, 2018).
One of the efforts towards more capable forecasters focuses on the foundation models (Bommasani et al., 2021), which address data-scarce scenarios by pre-training. More capable models support zero-shot forecasting, making inferences as fast as statistical methods and possessing large model capacity as deep models. Another aspect is to address uncertainty in time series forecasting. There is a growing research emphasis on probabilistic forecasting (Woo et al., 2024; Ansari et al., 2024). While parametric densities can be adopted as training objectives of probabilistic forecasting, they can be too specific to meet the heterogeneity of large-scale datasets, resulting in mode collapse in representation learning and over-smooth predictions (Figure 14-15). In this work, we introduce generative time series foundation models, which naturally address the uncertainty in forecasting.
@@ 2.2 Time Series Foundation Models
Recent research has concentrated on building versatile large time series models (Liang et al., 2024). With the advances made in large language models, Transformer has become the dominant architecture. Several works adapt Transformers to address the unique 2D-dimensionality and heterogeneity of time series (Woo et al., 2024; Liu et al., 2024a). Specifically, our work delves into tokenization and optimization. Models such as TimesFM (Das et al., 2023b), Timer (Liu et al., 2024a, b), and Time-MoE (Shi et al., 2024b) embed continuous values and fit unimodal distributions via MSE or quantile loss (Wen et al., 2017). However, prior loss may result in mode collapse because predictive distributions are highly divergent across different domains. Besides, these models cannot provide the confidence level of predictions, limiting reliability for decision-making. Based on continuous tokenization, Moirai (Woo et al., 2024) presents a probabilistic model learning a mixture of distributions, but this prior can still fail to accommodate complex distributions. Inspired by language modeling, Chronos (Ansari et al., 2024) discretizes series via bucket quantization, learning more flexible categorical distributions by cross-entropy. Still, discrete tokenization is applied at each time point, which can lead to long contexts. Also, the final performance can be sensitive to quantization techniques. Unlike before, we tame Transformers as native time series foundation models, learning flexible distributions without discrete tokenization.
@@ 2.3 Generative Modeling for Time Series
By addressing complicated distributions during pre-training, generative modeling has become a focal point in the development of various foundation models (Zhao et al., 2023; Liu et al., 2024c). While this direction for time series mostly focused on time series generation (Tashiro et al., 2021) and task-specific forecasters (Rasul et al., 2021; Shen & Kwok, 2023; Kollovieh et al., 2024), generative modeling for time series foundation models is hardly explored. With the comparable flexibility in distribution learning as language modeling, diffusion denoising (Sohl-Dickstein et al., 2015) and flow-matching (Lipman et al., 2022) have gained increasing prevalence in continuous-valued modalities (Lipman et al., 2024). Compared with diffusion denoising models, flow-matching provides a simple yet efficient framework. With fewer steps involved in the forward and reverse processes, large models based on flow-matching have shown superior performance in image generation (Esser et al., 2024).
Despite the connection in value continuity, generating images and future time series are fundamentally different tasks due to the autoregressive property of forecasting. Our proposed TimeFlow Loss is designed for autoregressive models to conduct conditional generation, which is a parameterized loss function (Zhang et al., 2018) for arbitrary distributions and enhances representation learning of foundation models.
@@ 3.2 Generative Models for Probabilistic Forecasting
Given a historical observation $x_{1:t}=\{x_{1},...,x_{t}\}$, the target of time series forecasting is to predict future time series $x_{t+1:t+f}=\{x_{t+1},\dots,x_{t+f}\}$. The task can be generally formulated as $p(x_{t+1:t+f}|\mathbf{h}_{t})$, where $\mathbf{h}_{t}=f_{\phi}(x_{1:t})$ is the learned representation from a deep model $f_{\phi}$. In probabilistic forecasting, explicit optimization objectives are utilized to predict the statistics of future series, e.g., MSE or quantile loss, which have specified $p$ as a prior distribution. While using one parametric density generally fits well on a small amount of data, it can be the major bottleneck for scaling time series foundation models. Inspired by the success of large generative models (Rombach et al., 2022; OpenAI, 2023; Esser et al., 2024), we introduce generative modeling to realize probabilistic forecasting:
@@
$g_{\theta}$ is a small trainable generative network conditioned on the learned representations of $f_{\phi}$, which is jointly optimized with $f_{\phi}$. While the generative model automatically fits the target distribution, it can sample raw predictions and calculating their statistics for probabilistic forecasting. The aim is conceptually related to conformal prediction (Vovk et al., 2005) but models uncertainty beyond prediction intervals.
@@ 4 Approach
In this work, we conduct a univariate pre-training paradigm, which adopts the S3 format proposed by Liu et al. (2024b) to address multivariate data. To mitigate value range discrepancy, we conduct normalization on time series individually per variable. Afterwards, we sample varying-length training samples with the maximum context length of $2880$. As a foundation model, Sundial is required to predict on out-of-distribution series with varied lengths during inference.
@@ 4.1 Sundial
As shown in Figure 2, the Sundial models consist of three parts: (1) time series tokenization, including a context-level re-normalization and a patch embedding that addresses any-length time series, (2) a Transformer backbone that learns the per-token representation of time series, and (3) TimeFlow Loss, a parameterized loss function to model the per-token distribution and generate raw series during inference. Intuitively, Sundial can be regarded as an ARMA (Auto-Regression and Moving-Average) deep model, i.e., Transformer learns token representations autoregressively. Conditioned on the lookback representations, TimeFlow transforms random noises into non-deterministic predictions.
@@ 4.1.1 Time Series Tokenization
We adopt stationarization (Liu et al., 2022), a non-parametric two-stage instance normalization conducted within each sample, which is initially proposed to mitigate non-stationarity of time series. Here, it helps to address temporal distribution shift and outlier ranges in input series, improving generalizability for zero-shot forecasting.
@@
where $\mathbf{h}_{i}\in\mathbb{R}^{D}$ and $D$ is the dimension of token embedding. Unlike point-level quantization (Ansari et al., 2024), we reserve original values without discrete quantization. It also reduces the context length (in token) of the Transformer.
@@ 4.1.2 Transformer Backbone
Given $N$ token embeddings $\{\mathbf{h}_{i}\}$, we adopt several crucial adaptations on a decoder-only Transformer to obtain per-token representations aggregated from all previous tokens. First, we adapt Pre-LN (Xiong et al., 2020) to improve pre-training stability. Second, we leverage a causal self-attention mechanism with RoPE (Su et al., 2024) that introduces the position information of patch tokens. It can be formulated as follows (the layer index is omitted for simplicity):